A clock showing correct time when at a rest loses 1 hour when it is moving.what is velocity

1
Expert's answer
2022-02-17T04:36:53-0500

Given:

"\Delta t=24\rm \: hr"

"\Delta \tau=23\rm \: hr"


The time delay equation says

"\Delta t=\frac{\Delta \tau}{\sqrt{1-v^2/c^2}}"

So, the required velocity

"v=c\sqrt{1-(\Delta \tau/\Delta t)^2}"

"v=c\sqrt{1-(23/24)^2}=0.29c"